OpenStudy (anonymous):

Apply the convolution theorem to find the inverse Laplace transform of the function: F(s) = s^2/(s^2+ 4)^2

OpenStudy (perl):

remind me what the convolution theorem (sounds convoluted)

OpenStudy (anonymous):

Laplace{f*g} = Laplace {f(t)}Laplace{g(t)} = F(s)G(s)

OpenStudy (anonymous):

s^2/(s^2+4) * (2/2)1/(s^2+4)

OpenStudy (anonymous):

is it familiar?

OpenStudy (anonymous):

from the original equation, I get \[s/(s^2+4) * s/(s^2+4)\]

OpenStudy (anonymous):

you have the tables?

OpenStudy (anonymous):

yes, both equations refer to cos equations

OpenStudy (anonymous):

ah you can do it that way too yes

OpenStudy (anonymous):

do you have a final answer? I think I may have one

OpenStudy (anonymous):

you are correct i think

OpenStudy (anonymous):

\[\int\limits_{0}^{t}\cos4xcos4(x-t)dx\] and so on

OpenStudy (anonymous):

sorry 2 there instead of 4

OpenStudy (anonymous):

yep I have that!